OpenStudy (anonymous):

Solve this word problem. Please help! A cosmetologist has a bottle of 7% hydrogen peroxide solution and a bottle of 4% hydrogen peroxide solution. The cosmetologist needs 300 milliliters of a 5% hydrogen peroxide solution for a hair dye. Write and solve a linear system to find how many milliliters of each solution the cosmetologist needs to mix together.

OpenStudy (anonymous):

@zepdrix

ILovePuppiesLol (ilovepuppieslol):

@imqwerty

jhonyy9 (jhonyy9):

7% mean 0,7 g in 100 mL solution 4% ---- 0,4 g in 100 mL solution the cosmotologist need 300 mL of 5% solution

jhonyy9 (jhonyy9):

do you know how you can making 5% solution fron 7% and 4% mixed ?

jhonyy9 (jhonyy9):

from - sorry

jhonyy9 (jhonyy9):

how many 7% and how many 4% will result 5% ?

jhonyy9 (jhonyy9):

for example 7+4=11/2=5,5% - not is right how will get 5% exactly ?

ILovePuppiesLol (ilovepuppieslol):

i have no idea but u seem like u know what u are doing @jhonyy9

jhonyy9 (jhonyy9):

@bookish610 check please my answers here - hope will help you